OpenStudy (anonymous):

algebra 2 question about functions. wait while I draw it

OpenStudy (anonymous):



OpenStudy (debbieg):

Do you mean \(f^{-1}(x)\)?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the problem has a positive 1.. is that weird?

OpenStudy (anonymous):

i know how to do the first 2 steps

OpenStudy (debbieg):

Yes, I'm not sure what that means. Someone yesterday posted the same kind of question. What have you covered in your class? Is this in a section on inverse functions??

OpenStudy (anonymous):

no, function notation

OpenStudy (anonymous):

there are a few inverse function questions too

OpenStudy (debbieg):

Where did your y go, and where did your 7 come from?

OpenStudy (anonymous):

i switched it

OpenStudy (anonymous):

wait woops

OpenStudy (anonymous):

its supposed to be 5x= y +12

OpenStudy (anonymous):



OpenStudy (anonymous):

you switch the x and the y variables

OpenStudy (debbieg):

OK, I think this is asking for in inverse function. Not sure though why your problem would use the "1" instead of "-1", that isn't correct notation. But let's assume it's asking you to find an inverse.

OpenStudy (anonymous):

ok

OpenStudy (debbieg):

Right, you are finding an inverse so you switch x and y. Now you said the function was: y=5x+12 So "switch the x and y" means to change the letter to y to x, and the letter x to y.

OpenStudy (debbieg):

It does NOT mean to just swap the positions of the terms, follow me? (that's what you did)

OpenStudy (anonymous):

ohhh

OpenStudy (debbieg):

So what do you get when you SWITCH x for y and y for x?

OpenStudy (anonymous):

so.. x= 5y+12?

OpenStudy (debbieg):

Right! Now solve for y!

OpenStudy (anonymous):

not sure how..

OpenStudy (anonymous):

divide both sides by 5?

OpenStudy (debbieg):

Once you do that, that is your inverse function, \(f^{-1}(x)\)

OpenStudy (debbieg):

Not yet, what do you need to do first? Before you divide?

OpenStudy (debbieg):

When working to solve an equation, always get all of the terms WITHOUT the variable on one side, and terms with the variable on the other, first. THEN you can divide to get the variable alone.

OpenStudy (anonymous):

so.. x-5y = 12?

OpenStudy (debbieg):

Dont move the 5y, move the 12. You are solving FOR y, so get the 5y term all alone.

OpenStudy (anonymous):

5y= x-12

OpenStudy (anonymous):

then divide by 5 on both sides

OpenStudy (debbieg):

Yes!

OpenStudy (anonymous):

y = x-12/ 5

OpenStudy (debbieg):

Oops, careful.... if what you really mean is \[y=\frac{ x-12 }{5 }\]then perfect. But what YOU WROTE is \[y=x-\frac{ 12 }{5 }\]which is wrong.

OpenStudy (anonymous):

thats what i mean

OpenStudy (anonymous):

the first way

OpenStudy (anonymous):

y = x-12 <-- this over 5

OpenStudy (debbieg):

OK, then next time just make sure you use () if you don't want to mess with the equation editorl y=(x-12)/5 would have been correct. :)

OpenStudy (anonymous):

ok

OpenStudy (debbieg):

Yes, you got it!